Solve. 2-3(x+5)=5(3-x)+8
Answer by persian52(161)

2-3(x+5)=5(3-x)+8
Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides.
-3(x+5)=-2+5(3-x)+8
Simplify the right-hand side of the equation.
-3(x+5)=-5x+21
Multiply -3 by each term inside the parentheses.
(-3x-15)=-5x+21
Remove the parentheses around the expression -3x-15.
-3x-15=-5x+21
Since -5x contains the variable to solve for, move it to the left-hand side of the equation by adding 5x to both sides.
-3x-15+5x=21
Since -3x and 5x are like terms, subtract 5x from -3x to get 2x.
2x-15=21
Since -15 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 15 to both sides.
2x=15+21
Add 21 to 15 to get 36.
2x=36
Divide each term in the equation by 2.
(2x)/(2)=(36)/(2)
Simplify the left-hand side of the equation by canceling the common terms.
x=(36)/(2)
Simplify the right-hand side of the equation by simplifying each term.
Answer: x=18